OpenStudy (fabgirl1994):

Given the following functions f(x) and g(x), solve (f ⋅ g)(2) and select the correct answer below: (1 point) f(x) = 2x2 − 10 g(x) = x + 9 A −22 B 22 C −66 D 66 2. Given the following functions f(x) and g(x), solve f over g (−4) and select the correct answer below: (1 point) f(x) = 4x − 4 g(x) = x − 1 A −20 B −4 C 4 D one fourth

OpenStudy (maddy1251):

First off, is f(x) = 2x2 − 10 written as \[2^{2} - 10\]

OpenStudy (maddy1251):

Oops

OpenStudy (fabgirl1994):

I understood the last one @Maddy1251 but this is worded different

OpenStudy (fabgirl1994):

Yes it is

OpenStudy (maddy1251):

\[2x^{2} - 10\]

OpenStudy (maddy1251):

Okay

OpenStudy (fabgirl1994):

Yes thats correct

OpenStudy (fabgirl1994):

The one you just posted

OpenStudy (maddy1251):

f(x) = 2x2 − 10 g(x) = x + 9 Now you know with how the last ones, we added them together?

OpenStudy (fabgirl1994):

Yes, so is it 2x2-10+x+9?

OpenStudy (maddy1251):

No. But when it said (f+g), we added them together. This time it says ( f x g) we we will multiply them together! Can you do that for me?

OpenStudy (maddy1251):

f(x) = 2x2 − 10 g(x) = x + 9 We multiply these two together

OpenStudy (fabgirl1994):

2x2-10x+9?

OpenStudy (maddy1251):

Well, f(x) = 2x^{2} - 10

OpenStudy (maddy1251):

(2x^2 -10)(x+9) Foil them for me. First, out, inner, last

OpenStudy (fabgirl1994):

Right

OpenStudy (mathmale):

To confirm: here we are MULTIPLYING f and g. You both seem to recognize that. Please show all steps of your multiplication.

OpenStudy (anonymous):

if fx Is (2x)2 that would be the answer

OpenStudy (mathmale):

(f * x)(x)= ( )( )

OpenStudy (maddy1251):

Yes what mathmale is getting as. If you put them into FOIL method you can then multiply all terms, then do whatever configuring you need to do.

OpenStudy (maddy1251):

f(x) = 2x2 − 10 g(x) = x + 9 so ; (2x^2-10)(x+9)

OpenStudy (fabgirl1994):

Okay that makes sense. So the answer is -22, or is there further steps?

OpenStudy (fabgirl1994):

Oh, I get it now. I forgot the parenthesis.

OpenStudy (maddy1251):

Did you multiply the two functions together?

OpenStudy (maddy1251):

I wanted to see what you got when you multiplied them together.

OpenStudy (fabgirl1994):

Let me do it real fast. Just a sec

OpenStudy (maddy1251):

Take your time. There is no rush

OpenStudy (fabgirl1994):

(2x^2-10)(x+9) 3x^2-19

OpenStudy (mathmale):

The work shared in your image is well done, and your answer (-22) is correct. whose work was that?

OpenStudy (maddy1251):

@mathmale had to have been hers. I will still on the step with her on multiplying them. :P

OpenStudy (mathmale):

OK. Good job (both of you).

OpenStudy (maddy1251):

Much thanks @mathmale , I have learned from watching the ones like you guys.

OpenStudy (fabgirl1994):

Can you help me with the second question, i posted on this question?

OpenStudy (maddy1251):

Yes! Let me go up and grab the equations

OpenStudy (fabgirl1994):

Okay I think I got the start of it right. I got

OpenStudy (maddy1251):

Given the following functions f(x) and g(x), solve f over g (−4) and select the correct answer below: (1 point) f(x) = 4x − 4 g(x) = x − 1 4x-4/x-1

OpenStudy (maddy1251):

Do they mean fractions, or dividing?

OpenStudy (fabgirl1994):

4x-4/x-1

OpenStudy (fabgirl1994):

Division

OpenStudy (maddy1251):

Okay! So 4x-4 divided by x-1

OpenStudy (fabgirl1994):

Would I divide 4x/x? So would it be 4x?

OpenStudy (fabgirl1994):

Then -4 because 4/-1?

OpenStudy (maddy1251):

Erm. Not quite. Hold on :)

OpenStudy (fabgirl1994):

Ok

OpenStudy (maddy1251):

 f(x) = 4x − 4 g(x) = x − 1

OpenStudy (maddy1251):



OpenStudy (fabgirl1994):

Correct

OpenStudy (maddy1251):

So it is 4

OpenStudy (fabgirl1994):

Okay, then what about the last part

OpenStudy (maddy1251):

Hmmm when it equals -4? If I remember that being correct

OpenStudy (fabgirl1994):

Yes

OpenStudy (fabgirl1994):

So the first part is 4x right?

OpenStudy (maddy1251):

I wonder if they mean Given the following functions f(x) and g(x), solve f over g (−4) and select the correct answer below: (1 point) f(x) = 4x − 4 g(x) = x − 1 4x-4 / x - 1 x= -4 4(-4)-4 / (-4) - 1 Is this what you were trying to get at from the beginning?

OpenStudy (maddy1251):

-16 - 4 / -5

OpenStudy (fabgirl1994):

Yes.

OpenStudy (maddy1251):

-20 / -5

OpenStudy (fabgirl1994):

So 4 is the answer?

OpenStudy (maddy1251):

-20 / - 5 = 4 huh well see when I divided the two, I got 4. So I had the answer from just dividing them like regular numbers, or you could do synthetical division but I didn't know if you knew how to do that.. Or, put them in fractional terms and substitute "-4" for x.

OpenStudy (maddy1251):

Yep, that's the answer. Do you understand how to get the answer?

OpenStudy (fabgirl1994):

Yes, I got it. Thanks so much for your help :)

OpenStudy (maddy1251):

Not a problemo :)

OpenStudy (maddy1251):

As long as you understand everything I am happy

OpenStudy (anonymous):

the answer is 4 even without replacing x by taking 4 as a common factor it will be 4(x-1)/x-1......x-1/x-1=1. so 4x1=4